# When AI Takes the Wheel: Security Analysis of Framework-Constrained Program Generation

### v1

@@ Abstract.
copyright: nonecopyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Large language models (LLMs) and low-code/no-code platforms are transforming software development. Popular LLMs (e.g., ChatGPT (OpenAI, 2024), Claude (Anthropic, 2023)) and low-code tools (e.g., Copilot Workspace (VSC, 2024), Cursor (cur, 2024)) enable faster prototyping and minimize the need for hand-written code. Even novice developers can now build complex software systems by providing description prompts, and then LLM-based systems “take the wheel” to handle technical implementation details. This evolution is happening at scale. According to a recent GitHub survey, 92% of developers are using AI coding tools (Shani and Staff, 2024). Google reports that over 25% of new code for their products is now generated by AI (Kelly, 2024), while Anthropic’s CEO predicts that AI will write 90% of code within 6 months and “essentially all of the code” within a year (Central, 2024). Although AI-assisted development improves productivity, it raises significant security concerns.
@@
Previous studies (Pearce et al., 2022; Perry et al., 2023) have examined security issues introduced by AI coding tools, and they demonstrated that AI-generated code often contains vulnerabilities. For example, Pearce et al. (Pearce et al., 2022) found that 40% of GitHub Copilot’s code suggestions contained potential vulnerabilities across various security-sensitive scenarios, such as SQL injection and path traversal. While recent research shows that newer LLM versions have improved, with vulnerability rates decreasing by 25% (Majdinasab et al., 2024), these studies mainly evaluate snippet-level security issues in isolated code completion tasks. However, LLMs now enable a ‘‘prompt-to-program’’ paradigm (a.k.a, vibe coding (vib, 2024)), where they complete all stages of software development from design to implementation. What remains unclear is how well these AI systems address security implications when they fully “take the wheel”, influencing architectural decisions and system-wide security considerations across complete applications.
@@
Security risks in modern software systems become especially challenging for framework-constrained applications that must adhere to specific security models and practices. For instance, Android applications require solid permission setups, secure inter-component communication, and careful storage of sensitive data across the system (Faruki et al., 2014; Liu et al., 2022). Chrome extensions pose similar challenges, catering to strict Content Security Policies (CSP), managing privileged API access, and ensuring secure communication between isolated components (Fass et al., 2021). In classical software development, experienced developers typically understand these framework constraints and their security implications. However, as LLMs step into the entire development process, developers may have less direct involvement in these critical security decisions. There is limited understanding of how effectively these AI systems address complex security requirements across different framework-constrained environments when they generate complete software systems based on high-level requirements.
@@ 2.2. Security Concerns in AI Solutions
While AI coding tools significantly improve developer productivity, they also introduce security concerns that undermine their reliability in software development. Previous studies (Pearce et al., 2022; Hamer et al., 2024; Fu et al., 2025; She et al., 2023) have highlighted that AI coding tools like GitHub Copilot potentially recommend insecure code, and Pearce et al. (Pearce et al., 2022) found that 40% of the 1,689 code snippets generated by GitHub Copilot across various programming scenarios contained security vulnerabilities. These findings extend to real-world development scenarios. Fu et al. (Fu et al., 2025) analyzed GitHub Projects using AI tools like GitHub Copilot and CodeWhisperer, revealing that 30% of the generated code snippets contained security weaknesses, with 43 different Common Weakness Enumeration (CWE) vulnerabilities identified in real-world projects. Besides that, Niu et al. (Niu et al., 2023) demonstrated that AI coding tools could inadvertently leak private information, and were susceptible to membership inference attacks. Additionally, prior studies (Liu et al., 2024; Yu et al., 2024) had shown that AI tools like ChatGPT can produce hallucinations and quality issues in program generation. Moreover, Perry et al. (Perry et al., 2023) conducted a user study revealing a more concerning issue: Developers with access to AI coding tools not only produced more insecure code than those without such tools, but also exhibited greater overconfidence in the security of their solutions (i.e., frequently rating insecure AI solutions as secure). While prior research effectively highlights these snippet-level security issues in code completion tasks, LLMs have significantly advanced in their coding abilities to generate complete complex programs. Developers increasingly rely on these tools to develop complex software systems, sometimes generating entire programs or applications with a minimal level of human involvement. Thus, a critical question remains unanswered: whether the implications about security vulnerabilities in AI-generated code snippets persist (or are amplified) in complete AI-generated software systems.
@@ 3.1. Problem Definition
We focus on evaluating potential security vulnerabilities in framework-constrained Chrome extensions generated by LLMs. Unlike previous studies that mainly examined self-contained software tasks or isolated code snippets (Pearce et al., 2022; Majdinasab et al., 2024), our investigation analyzes complete extension implementations that must correctly navigate Chrome’s “isolated worlds” security model across different privilege boundaries. This requires analyzing complex interactions between extension components operating at different privilege levels, including content scripts, background scripts, and web pages, as well as their communication channels. We do not attempt to demonstrate concrete exploits or verify that LLMs implement the correct functionalities as described, but rather focus on identifying vulnerable data flows and architectural weaknesses that could lead to privilege escalation, data exfiltration, and other extension-specific security threats within this framework-constrained environment.
@@ 3.2. Evaluating Extensions Using Static Analysis
We employ CoCo (Yu et al., 2023) as our primary analysis tool. CoCo is a static analysis framework that performs coverage-guided concurrent abstract interpretation to detect vulnerabilities in browser extensions. Using CoCo, we analyze the generated extensions through abstract interpretation with concurrent taint propagation, tracking data flows across multiple execution paths. Our analysis focuses on identifying potentially dangerous data flows that could lead to security vulnerabilities. We specifically track flows between security-sensitive sources (e.g., external message-passing APIs) and potentially dangerous sinks (e.g., file downloads, storage access). Using CoCo’s concurrent analysis capabilities, we efficiently detect vulnerabilities that arise from complex interactions between content scripts, background scripts, and web pages.
@@ 3.3. Scenario Design and Implementation
We design a structured approach to perform a systematic empirical evaluation of potential security risks in LLM-generated Chrome extensions. We construct a prompt dataset that instructs LLMs to generate extensions with various functionalities. Drawing from three prior research studies (Fass et al., 2021; Yu et al., 2023; Somé, 2019) that identified and released vulnerable Chrome extensions, we leverage their documented vulnerability patterns and lists of extensions that exhibited security weaknesses. For each vulnerable extension, we create a natural language prompt that describes its core functionality, forming our ChromeSecBench dataset. We provided these prompts to popular LLMs (e.g., GPT-4o, Claude-3.5-Sonnet) to generate complete extension implementations and analyzed the security properties of the generated code. Our overall experimental workflow is depicted in Figure 3.
@@
In step  1, for each identified vulnerable extension, we write a number of natural language prompts that describe the extension’s intended functionalities. These prompts contain descriptions of the desired Chrome extension’s core features and behaviors that will be used to instruct LLMs to generate complete implementations (like Figure 1 shows how a developer would use AI assistance in practice). To ensure prompt quality and consistency, we used ChatGPT (GPT-4o) as an editorial assistant. It was tasked with reviewing our manually-written description alongside the original extension’s source code to help refine grammar, remove ambiguity, and suggest any missing functional elements. For example, it helped refine a general prompt like “Chrome extension that automates job applications…” into the more functionally complete version used in our study: “Chrome extension that automates job applications through Cutback.io…” by pointing out the missing target website. Note that these prompts remained human-authored and focused on high-level functionality without implementation details, mirroring how end-users or non-expert developers typically specify requirements. While our system prompt contains general security guidance (”Ensure the extension works correctly and securely”), we deliberately avoided specific security constraints or technical implementation details (e.g., validate all external URLs before fetching data). This design choice allows us to evaluate LLMs’ intrinsic understanding of framework security models under typical usage conditions, rather than testing the effectiveness of detailed security prompt engineering. For consistency, we restrict the original manifest file specifications in the input prompts to essential permissions and basic configurations. In developing these scenarios, we rely on vulnerable extensions identified in  (Fass et al., 2021; Yu et al., 2023; Somé, 2019) that include complete source code and documentation. We exclude extensions without available source code or those exceeding 1,000 lines of code to maintain manageable complexity and relevance. This process results in our ChromeSecBench dataset  2, which comprises distinct scenarios for 140 Chrome extensions.
@@
Next, in  3, we feed the prompt to popular LLMs to generate complete Chrome extension implementations. In our study, we evaluate advanced models like GPT-4o and Claude-3.5-Sonnet, as well as reasoning-focused models like o3-mini and DeepSeek-r1. Each model is prompted up to 12 times per prompt to account for potential variations in the outputs. The complete responses are processed to extract and separate distinct components (HTML, JavaScript, and CSS files), which are then combined with the original manifest specifications to create functional Chrome extensions in  4. During this process, some generated implementations are discarded in  4a when the static analysis tool fails to process them, often due to critical syntax errors or non-standard code structures that exceed the tool’s capabilities.
@@
Finally, in  5, we evaluate each generated Chrome extension for security vulnerabilities. This evaluation is mainly conducted using CoCo  5a, which employs coverage-guided concurrent abstract interpretation to perform taint analysis and data flow tracking. We choose CoCo due to its proven strong precision ( 94.3%) and recall ( 99.2%) in vulnerability detection for Chrome extensions over previous tools (Yu et al., 2023), especially suitable for our analyzed samples’ characteristics: small size, no obfuscation, and simple constructs that allow accurate automated analysis. In addition, we perform a manual code review  5b on a random sample to verify their data flow behaviors and confirm CoCo performance in our specific context. Our evaluation focuses specifically on the sensitive API endpoints and data flow patterns listed in the original CoCo framework (Yu et al., 2023). We do not assess whether these vulnerabilities are exploitable in real-world attack scenarios or whether an attacker could successfully reach and exploit the identified security flaws. This scope limitation is deliberate, as our primary objective is to understand whether and how LLMs naturally introduce potentially dangerous data flow patterns during program generation, rather than proving the practical exploitability of each identified vulnerability. The analysis results are then collected and synthesized in  7 to provide comprehensive insights into the security characteristics of LLM-generated programs.
@@ 3.4. Experimental Setup
We ran all experiments on Ubuntu 20.04.3 LTS with an AMD Ryzen 9 5950X 16-Core Processor (3.4 GHz) and 64GB RAM. While we completed steps  1 and  2 manually, we developed automated Python scripts to handle steps  3, 4a, and  5. All scripts are implemented in Python 3.8.5, using official API interfaces to interact with the language models: OpenAI API for GPT-4o and o3-mini (with the default reasoning effort set to medium), Anthropic API for Claude-3.5-Sonnet (version 2024-10-22), Google Vertex AI API for Gemini-1.5-Pro, Gemini-2-Pro-exp-02-05 and Gemini-2-Flash-Thinking-exp-01-21, DeepSeek API for DeepSeek-V3 and DeepSeek-R1, and Google Vertex AI for accessing the Llama-3.1-405B model. We set temperature to 1.0 to get diverse implementations and max_tokens to 10,000 so the models could generate complete extensions. For step  5, we employ the open-source CoCo framework (Yu et al., 2023) with its default configuration settings for vulnerability analysis. The complete experimental results, generated extensions, analysis artifacts, and all associated prompts (including the system prompts for generation and the prompts used to refine our written descriptions) are available in our public repository at https://github.com/yueyueL/ChromeSecBench.
@@ 4. Experimental Analysis
Figure 4. Vulnerability Occurrence Across Different Extension Scenarios (darker colors indicate higher percentage of vulnerable generations per scenario)
To systematically analyze AI-generated Chrome extensions, we developed ChromeSecBench, comprising 140 natural language scenarios based on real-world vulnerable extensions. To account for the inherent variability in LLM outputs, we prompt each model 12 times per scenario, resulting in 1,680 generation attempts per model. We analyze the security implications across three key dimensions: diversity of scenarios, diversity of LLMs, and diversity of vulnerability types, to understand how these factors contribute to security risks when AI ”takes the wheel” in framework-constrained program generation.
Static Analysis Validation.  To further verify the performance of vulnerability detection results from CoCo, we randomly selected 100 samples selected from each set labeled vulnerable and non-vulnerable, and manually audited them. The first author manually examined all 200 samples, with results verified by the third author to ensure accuracy and eliminate bias. Our manual audit confirmed zero false positives in the ”vulnerable” set, affirming CoCo’s high precision in our context. We identified two false negatives, one resulting from an undefined sink rule for XMLHttpRequest.open and the other from a complex data flow pattern with indirect API calls. These types of limitations are consistent with known challenges in static analysis (Yu et al., 2023). Overall, this validation demonstrates CoCo’s high accuracy (98% recall, 99% F1 score) for our specific evaluation context, where the extensions’ small size ( 200 lines), lack of obfuscation, and straightforward programming constructs align well with CoCo’s detection capabilities. This high accuracy provides confidence that our subsequent vulnerability analysis accurately reflects the security properties of LLM-generated Chrome extensions.

### v2

@@ Abstract.
Large Language Models, Program Generation, Chrome Extensions, Security Vulnerabilities
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773142isbn: 979-8-4007-2025-3/2026/04ccs: Security and privacy Software and application security
@@ 1. Introduction
Large language models (LLMs) and low-code/no-code platforms are transforming software development. Popular LLMs (e.g., ChatGPT (OpenAI, 2024), Claude (Anthropic, 2023)) and low-code tools (e.g., Copilot Workspace (cop, 2024), Cursor (cur, 2024)) enable faster prototyping and minimize the need for hand-written code. Even novice developers can now build complex software systems by providing description prompts, and then LLM-based systems “take the wheel” to handle technical implementation details. This evolution is happening at scale. According to a recent GitHub survey, 92% of developers are using AI coding tools (Shani and Staff, 2024). Google reports that over 25% of new code for their products is now generated by AI (Kelly, 2024), while Anthropic’s CEO predicts that AI will write 90% of code within 6 months and “essentially all of the code” within a year (Central, 2024). Although AI-assisted development improves productivity, it raises significant security concerns.
@@
Previous studies (Pearce et al., 2022; Perry et al., 2023) have examined security issues introduced by AI coding tools, and they demonstrated that AI-generated code often contains vulnerabilities. For example, Pearce et al. (Pearce et al., 2022) found that 40% of GitHub Copilot’s code suggestions contained potential vulnerabilities across various security-sensitive scenarios, such as SQL injection and path traversal. While recent research shows that newer LLM versions have improved, with vulnerability rates decreasing by 25% (Majdinasab et al., 2024), these studies mainly evaluate snippet-level security issues in isolated code completion tasks. However, LLMs now enable a ‘‘prompt-to-program’’ paradigm (a.k.a. vibe coding (vib, 2024)), where they complete all stages of software development from design to implementation. What remains unclear is how well these AI systems address security implications when they fully “take the wheel”, influencing architectural decisions and system-wide security considerations across complete applications.
@@
Security risks in modern software systems become especially challenging for framework-constrained applications that must adhere to specific security models and practices. For instance, Android applications require solid permission setups, secure inter-component communication, and careful storage of sensitive data across the system (Faruki et al., 2014; Liu et al., 2022). Chrome extensions pose similar challenges, adhering to strict Content Security Policies (CSP), managing privileged API access, and ensuring secure communication between isolated components (Fass et al., 2021). In classical software development, experienced developers typically understand these framework constraints and their security implications. However, as LLMs step into the entire development process, developers may have less direct involvement in these critical security decisions. There is limited understanding of how effectively these AI systems address complex security requirements across different framework-constrained environments when they generate complete software systems based on high-level requirements.
@@ 2.2. Security Concerns in AI Solutions
While AI coding tools significantly improve developer productivity, they also introduce security concerns that undermine their reliability in software development. Previous studies (Pearce et al., 2022; Hamer et al., 2024; Fu et al., 2025; She et al., 2023) have highlighted that AI coding tools like GitHub Copilot potentially recommend insecure code, and Pearce et al. (Pearce et al., 2022) found that 40% of the 1,689 code snippets generated by GitHub Copilot across various programming scenarios contained security vulnerabilities. These findings extend to real-world development scenarios. Fu et al. (Fu et al., 2025) analyzed GitHub Projects using AI tools like GitHub Copilot and CodeWhisperer, revealing that 30% of the generated code snippets contained security weaknesses, with 43 different Common Weakness Enumeration (CWE) vulnerabilities identified in real-world projects. Besides that, Niu et al. (Niu et al., 2023) demonstrated that AI coding tools could inadvertently leak private information, and were susceptible to membership inference attacks. Additionally, prior studies (Liu et al., 2024; Yu et al., 2024) have shown that AI tools like ChatGPT can produce hallucinations and quality issues in program generation. Moreover, Perry et al. (Perry et al., 2023) conducted a user study revealing a more concerning issue: Developers with access to AI coding tools not only produced more insecure code than those without such tools, but also exhibited greater overconfidence in the security of their solutions (i.e., frequently rating insecure AI solutions as secure). While prior research effectively highlights these snippet-level security issues in code completion tasks, LLMs have significantly advanced in their coding abilities to generate complete complex programs. Developers increasingly rely on these tools to develop complex software systems, sometimes generating entire programs or applications with a minimal level of human involvement. Thus, a critical question remains unanswered: whether the implications of security vulnerabilities in AI-generated code snippets persist (or are amplified) in complete AI-generated software systems.
@@ 3.1. Problem Definition
We focus on evaluating potential security vulnerabilities in framework-constrained Chrome extensions generated by LLMs. Unlike previous studies that mainly examined self-contained software tasks or isolated code snippets (Pearce et al., 2022; Majdinasab et al., 2024), our study analyzes complete extension implementations that must correctly navigate Chrome’s “isolated worlds” security model across different privilege boundaries. This requires analyzing complex interactions between extension components operating at different privilege levels, including content scripts, background scripts, and web pages, as well as their communication channels. We do not attempt to demonstrate actual exploits or check whether the extensions function correctly. Instead, we focus on identifying vulnerable data flows that could allow privilege escalation, data exfiltration, and other extension-specific attacks.
@@ 3.2. Evaluating Extensions Using Static Analysis
We employ CoCo (Yu et al., 2023) for our analysis. CoCo is a static analysis framework that performs coverage-guided concurrent abstract interpretation to detect vulnerabilities in browser extensions. Using CoCo, we analyze the generated extensions through abstract interpretation with concurrent taint propagation, tracking data flows across multiple execution paths. Our analysis focuses on identifying potentially dangerous data flows that could lead to security vulnerabilities. We specifically track flows between security-sensitive sources (e.g., external message-passing APIs) and potentially dangerous sinks (e.g., file downloads, storage access). Using CoCo’s concurrent analysis capabilities, we efficiently detect vulnerabilities that arise from complex interactions between content scripts, background scripts, and web pages.
@@ 3.3. Scenario Design and Implementation
We design a structured approach to perform a systematic empirical evaluation of potential security risks in LLM-generated Chrome extensions. We construct a prompt dataset that instructs LLMs to generate extensions with various functionalities. Drawing from three prior research studies (Fass et al., 2021; Yu et al., 2023; Somé, 2019) that identified and released vulnerable Chrome extensions, we leverage their documented vulnerability patterns and lists of extensions that exhibited security weaknesses. For each vulnerable extension, we create a natural language prompt that describes its core functionality, forming our ChromeSecBench dataset. We provide these prompts to popular LLMs (e.g., GPT-4o, Claude-3.5-Sonnet) to generate complete extension implementations and analyze the security properties of the generated code. Our overall experimental workflow is depicted in Figure 3.
@@
In step  1, for each identified vulnerable extension, we write a number of natural language prompts that describe the extension’s intended functionalities. These prompts contain descriptions of the desired Chrome extension’s core features and behaviors that will be used to instruct LLMs to generate complete implementations (like Figure 1 shows how a developer would use AI assistance in practice). To ensure prompt quality and consistency, we used ChatGPT (GPT-4o) as an editorial assistant. It was tasked with reviewing our manually-written description alongside the original extension’s source code to help refine grammar, remove ambiguity, and suggest any missing functional elements. For example, it helped refine a general prompt like “Chrome extension that automates job applications…” into the more functionally complete version used in our study: “Chrome extension that automates job applications through Cutback.io…” by pointing out the missing target website. Note that these prompts remained human-authored and focused on high-level functionality without implementation details, the way developers describe their needs when seeking AI assistance for coding tasks. Our system prompt includes general security guidance (i.e., “Ensure the extension works correctly and securely”), but we avoid specific security constraints or implementation details (e.g., “validate all external URLs before fetching data”). This design lets us assess how well LLMs grasp framework security on their own in everyday use, instead of measuring the impact of detailed security prompts. For consistency, we restrict the original manifest file specifications in the input prompts to essential permissions and basic configurations. In developing these scenarios, we rely on vulnerable extensions identified in (Fass et al., 2021; Yu et al., 2023; Somé, 2019) that include complete source code and documentation. We exclude extensions if they have no source code or are over 1,000 lines of code. This helps keep the complexity manageable and ensure relevance. This yields our ChromeSecBench dataset  2, which comprises distinct scenarios for 140 Chrome extensions.
@@
Next, in  3, we feed the prompt to popular LLMs to generate complete Chrome extension implementations. In this study, we evaluate advanced models (e.g., GPT-4o and Claude-3.5-Sonnet), as well as reasoning-focused models (e.g., o3-mini and DeepSeek-R1). To account for potential variations in the output, we run each prompt up to 12 times on each model. The complete responses are processed to extract and separate distinct components (HTML, JavaScript, and CSS files), which are then combined with the original manifest specifications to create functional Chrome extensions in  4. We discard generated implementations (in 4a) that the static analysis tool cannot process due to critical syntax errors or non-standard code structures.
@@
Finally, in  5, we analyze the generated extensions for security vulnerabilities. This evaluation is mainly conducted using CoCo  5a, which employs coverage-guided concurrent abstract interpretation to perform taint analysis and data flow tracking. We choose CoCo due to its proven strong precision ( 94.3%) and recall ( 99.2%) in vulnerability detection for Chrome extensions over previous tools (Yu et al., 2023), especially suitable for our analyzed samples’ characteristics: small size, no obfuscation, and simple constructs that allow accurate automated analysis. In addition, we perform a manual code review  5b on a random sample to verify their data flow behaviors and confirm CoCo performance in our specific context. Our evaluation focuses specifically on the sensitive API endpoints and data flow patterns listed in (Yu et al., 2023). We do not assess if these vulnerabilities are exploitable in real-world attack scenarios. We take this approach because our main goal is to understand whether LLMs inherently introduce potentially dangerous data flow patterns during program generation, rather than demonstrating the practical exploitability of each identified vulnerability. The results are aggregated in  7 to provide insights into the security properties of LLM-generated programs.
@@ 3.4. Experimental Setup
We ran all experiments on Ubuntu 20.04.3 LTS with an AMD Ryzen 9 5950X 16-Core Processor (3.4 GHz) and 64GB RAM. While we completed steps  1 and  2 manually, we developed automated Python scripts to handle steps  3, 4a, and  5. We wrote all scripts in Python 3.8.5. To access the models, we used their official APIs: OpenAI API for GPT-4o and o3-mini (with the default reasoning effort set to medium), Anthropic API for Claude-3.5-Sonnet (version 2024-10-22), Google Vertex AI API for Gemini-1.5-Pro, Gemini-2-Pro-exp-02-05 and Gemini-2-Flash-Thinking-exp-01-21, DeepSeek API for DeepSeek-V3 and DeepSeek-R1, and Google Vertex AI for accessing the Llama-3.1-405B model. We set the temperature to 1.0 to get diverse implementations and max_tokens to 10,000 so the models could generate complete extensions. For step  5, we employ the open-source CoCo framework (Yu et al., 2023) with its default configuration settings for vulnerability analysis. The complete experimental results, generated extensions, analysis artifacts, and all associated prompts (including the system prompts for generation and the prompts used to refine our written descriptions) are available in our public repository at https://github.com/yueyueL/ChromeSecBench.
@@ 4. Experimental Analysis
Figure 4. Vulnerability Occurrence Across Different Extension Scenarios (darker colors indicate a higher percentage of vulnerable generations per scenario)
To thoroughly analyze AI-generated Chrome extensions, we developed ChromeSecBench, comprising 140 natural language scenarios based on real-world vulnerable extensions. To account for the inherent variability in LLM outputs, we prompt each model 12 times per scenario, resulting in 1,680 generation attempts per model. Our analysis considers three dimensions: diversity of scenarios, diversity of LLMs, and diversity of vulnerability types, to understand how these factors contribute to security risks when AI ”takes the wheel” in framework-constrained program generation.
Static Analysis Validation.  To further verify the performance of vulnerability detection results from CoCo, we randomly selected 100 samples from each set labeled vulnerable and non-vulnerable, and manually audited them. The first author manually examined all 200 samples, with results verified by the third author to ensure accuracy and eliminate bias. We did not find any false positives. We identified two false negatives, one resulting from an undefined sink rule for XMLHttpRequest.open and the other from a complex data flow pattern with indirect API calls. These types of limitations are consistent with known challenges in static analysis (Yu et al., 2023). Overall, CoCo achieves high accuracy (98% recall, 99% F1 score) for our specific evaluation context. The performance is expected since the extensions are small ( 200 lines), not obfuscated, and use simple code patterns that CoCo handles well. Thus, this high accuracy provides confidence that our subsequent vulnerability analysis accurately reflects the security properties of LLM-generated Chrome extensions.